* Assume a Class C network with the ID of 192.168.12.0 is to be subnetted into 8 smaller networks, complete the following:

1. How many host bits will you need to borrow to use for network information?

2. How many hosts will each subnet be able to support?

3. What is the CIDR notation for each network?

4. What is the subnet mask for each network?

5. What is the broadcast address of the subnet with the extended network prefix of 192.168.12.0?

6. What is the extended network prefix of the 4th subnet?

7. What is the first usable host address in the 5th subnet?

8. What is the last usable host address in the 5th subnet?

9. Express the last octet of the subnet mask in binary:

10. Which subnet would the host whose last octet is 00101111 reside?
Answer #1

******************************************************************************************
Please Upvote the answer as it matters to me a lot :)
*****************************************************************************************
As per HomeworkLib expert answering guidelines,Experts are supposed to answer only certain number of questions/sub-parts in a post.Please raise the remaining as a new question as per HomeworkLib guidelines.
******************************************************************************************

since it is given that the network is in class C
so there will be 24 bits in network part and 8 bits in the host part as per class C definition
since we need to divide the network into 8 smaller networks

1)we need 3 bits to represent the 8 smaller networks
so we need to borrows 3 bits from the host so that host will only have
5 bits now

2)since now each subnet has 5 bits, we can represent at max
2^32 hosts in a single subnet

3)
the first 3 bits in the last octet are the subnet bits
192.168.12.00000000 will be the first network
= 192.168.12.0/27 (24+3)

192.168.12.00100000 will be the 2nd network
= 192.168.12.32/27 (24+3)
192.168.12.01000000 will be the 3rd network
= 192.168.12.64/27 (24+3)
192.168.12.01100000 will be the 4th network
= 192.168.12.96/27 (24+3)
192.168.12.10000000 will be the 5th network
= 192.168.12.128/27 (24+3)
192.168.12.10100000 will be the 6th network
= 192.168.12.160/27 (24+3)
192.168.12.11000000 will be the 7th network
= 192.168.12.192/27 (24+3)
192.168.12.11100000 will be the 8th network
= 192.168.12.224/27 (24+3)

4)
the subnet mask for the networks is the same which is
first 27 bits 1's and rest all 0's =
11111111.11111111.11111111.11100000
= 255.255.255.224